# Chiral Gapless Spin Liquid in Hyperbolic Space

### v2

@@ Abstract
We analyze the Kitaev model on the $\{9,3\}$ hyperbolic lattice. The $\{9,3\}$ is formed by a regular 3-coordinated tiling of nonagons, where the 3-color coding of bonds according to the inequivalent Kitaev Ising spin couplings yields the natural generalization of the original Kitaev model for Euclidean regular honeycomb tiling. Upon investigation of the bulk spectrum for large finite size droplets, we identify a gapless chiral $\mathbb{Z}_{2}$ spin liquid state featuring spontaneous time reversal symmetry breaking. Due to its non-commutative translation group structure, such type of hyperbolic spin liquid is conjectured to feature chiral quasiparticles with a potentially non-Abelian Bloch profile.
@@
Figure 1: (a) Droplet of the $\{9,3\}$ hyperbolic lattice with Kitaev-colored bonds for each coupling $J_{x},J_{y},J_{z}$. The inset shows a schematic of the four Majorana fermions $\hat{b}_{j}^{\alpha},\hat{c}_{j}$ introduced at site $j$. (b) Definition of the positive direction for each link $\hat{u}_{jk}$. If all corresponding eigenvalues $u_{jk}$ are $+1$ ($-1$), the vortex-free (vortex-filled) sector is realized. (c) Schematic phase diagram of the $\{9,3\}$ Kitaev model in the $J_{x}+J_{y}+J_{z}=1$ plane. If one of the coupling constants is larger than the sum of the other two, the Majorana Hamiltonian is gapped. Otherwise we find a gapless, chiral phase.
@@ Introduction.
Quantum spin liquids have been in the focus of contemporary research on strongly correlated electron systems for five decades [1]. In experiment, the unambiguous identification of such an intricate quantum fluid state of matter is notoriously difficult, and rarely, if ever, possible beyond reasonable doubt [2]. In theory, our understanding has progressed at various frontiers, yet at highly different rates. Due to the seminal work by Wen [3], an assumingly exhaustive classification of quantum spin liquids is given from their associated emergent gauge structure at low energies. There usually is, however, still a long way to go from quantum order classifications, and their related parton mean fields, to a refined microscopic realization and stabilization of such states. This is why, in particular for quantum spin liquids and quantum many-body phases in general, it is highly desirable to accomplish parent Hamiltonians whose ground state unambiguously matches such a state. Typically, a combination of optimization ansatz and educated guess is employed to scan a multi-dimensional Hamiltonian parameter space [4, 5, 6].
@@
The Kitaev paradigm of spin liquid parent Hamiltonians has proven particularly successful [7]. Conceived as a bond-dependent anisotropic Ising Hamiltonian between neighboring sites on the honeycomb lattice, it has brought about various exactly solvable instances of quantum spin liquids with an emergent $\mathbb{Z}_{2}$ gauge structure. In terms of spin liquids featuring topological order, it includes the gapped Abelian toric code liquid and, upon consideration of an external magnetic field [7] or through spontaneous time-reversal (T) symmetry breaking via lattice graphs with odd loops [8, 9], a gapped non-Abelian chiral Majorana liquid with spinon quasiparticles which resemble Pfaffian liquid quasiparticles [10, 11]. The Kitaev model further harbors the gapless $\mathbb{Z}_{2}$ spin liquid centered around isotropic Ising couplings. Further progress has been reached by Hermanns and Trebst [12] who first employed that if a unique bond attribution to three subsets can be made for a semi-regular tricoordinated tiling, this still yields exact solvability of the Kitaev model. Upon transcending from two to three spatial dimensions, this enabled them to define Kitaev spin liquid states with a (potentially unstable [13]) spinon Fermi surface as well as chiral spin liquid states with crystalline gauge ordering of $\mathbb{Z}_{2}$ fluxes [14].
@@
Since the nature of the lattice strongly affects the magnetic frustration profile, any parent Hamiltonian search for quantum spin liquids seeks to consider all variations of potentially interesting lattices. Conventionally, this implies to be constrained to tessellations of Euclidean flat space. Recent progress on hyperbolic matter, however, has broadened the view to also consider Hamiltonians operating on negatively curved spaces and lattice tessellations thereof [15]. While the field initially started from unconventional non-interacting states and their description in terms of symmetry protected topological phases [16, 17, 18] and hyperbolic band theory [19, 20, 21, 22], it is now starting to reach the realms of quantum many-body physics through metal-insulator transitions, charge density wave states, and magnetism [23, 24], fractional Chern insulators [25], and lattice-regularized formulations of AdS-CFT correspondence [26, 27]. An immediate platform for tabletop experimental realizations of hyperbolic space presents itself to be quantum and classical circuit networks [28, 29, 30].
@@
In this Letter, we introduce the notion of two-dimensional (2D) hyperbolic spin liquids in the framework of a generalized Kitaev model. Applying the same logical thread as in Euclidean space, the focus of consideration lies on tricoordinated lattice tessellations in the hope of reaching an exactly solvable Hamiltonian. While in principle any of the vast hyperbolic space tessellations that fulfill this tricoordination deserves further investigation, we will particularize to a microscopic model for an exotic spin liquid state that features gapless excitations, and yet spontaneously breaks time reversal symmetry. It has no immediate analogue in Euclidean space, and hence expands the hitherto known landscape of spin liquids to uncharted territory.
@@
Realizing gapless excitations for a Kitaev Hamiltonian is intimately tied to regular tiling, i.e., tessellations with only one type of regular polygon, while spontaneous T symmetry breaking requires odd-sided loops. Such a tessellation of 2D Euclidean space by regular odd-sided tilings does not exist. Instead, one would have to resort to semi-regular tilings such as the Fisher star lattice [8, 31] or non-Archimedean realizations [9] such as the pentaheptite lattice both of which feature gapped chiral Kitaev spin liquids. This motivates us to transcend to 2D hyperbolic space to realize tricoordinated, regular odd-loop tilings. Among the vast tessellations that fulfill the constraint, the $\{9,3\}$ lattice Fig. 1(a) stands out, as it features the smallest loop size of hyperbolic tiling whose bond coloring appears highly canonical, and even suggests itself to be the most natural generalization of the Euclidean honeycomb Kitaev model. Our analysis of the $\{9,3\}$ hyperbolic Kitaev model finds a gapless state which, due to the presence of odd loops, features spontaneous time reversal symmetry breaking. We thus show the Kitaev model to harbor a gapless, chiral $\mathbb{Z}_{2}$ spin liquid state whose nature of excitations, however, is conjectured to be intimately dependent on the non-commutative translation group inherent to hyperbolic space.
@@
$\hat{H}=-J_{x}\sum_{\langle i,j\rangle_{x}}\hat{\sigma}_{i}^{x}\hat{\sigma}_{j}^{x}-J_{y}\sum_{\langle i,j\rangle_{y}}\hat{\sigma}_{i}^{y}\hat{\sigma}_{j}^{y}-J_{z}\sum_{\langle i,j\rangle_{z}}\hat{\sigma}_{i}^{z}\hat{\sigma}_{j}^{z}$ (1)
on a three-colored $\{9,3\}$ lattice. There, $\hat{\sigma}_{i}^{\alpha}$ denote Pauli matrices acting on site $i$, and $J_{\alpha}>0$ are the coupling constants along the respective link of type $\alpha=x,y,z$. The different link types are shown in Fig. 1(a) as red, green and blue lines, respectively. Although model (1) can be solved exactly for all tricoordinated three-colorable hyperbolic lattices, the $\{9,3\}$ appears to be the natural hyperbolic extension of the original Euclidean model due to the regular coloring scheme of repeated $x$-, $y$-, and $z$-links upon going counterclockwise around each plaquette (Fig. 1). Invoking $(p-2)(q-2)>4$ obeyed by all hyperbolic tilings, this canonical coloring scheme exists for all $\{p,3\}$ lattices with $p\equiv 0\mod 3,\;p>6$, so that the $\{9,3\}$ lattice is the one with smallest loop size.
The Hamiltonian (1) possesses a set of conserved quantities, one associated with each nonagon plaquette, given by the mutually commuting Wilson loop operators
$\hat{W}_{p}=\hat{K}_{12}\hat{K}_{23}\hat{K}_{34}\hat{K}_{45}\hat{K}_{56}\hat{K}_{67}\hat{K}_{78}\hat{K}_{89}\hat{K}_{91},$ (2)
where $\hat{K}_{ij}=\hat{\sigma}_{i}^{\alpha}\hat{\sigma}_{j}^{\alpha}$, and the sites $i$ and $j$ are labeled by indices $i,j=1,\cdots,9$ counterclockwise around a plaquette and connected by a bond of type $\alpha$. Since all plaquette operators commute, the Hamiltonian attains a block diagonal form according to the $\hat{W}_{p}$ eigenvalues. In contrast to Kitaev models with even-loop plaquettes, these are given by ±$\pm\mathrm{i}$ rather than ±$\pm 1$, implying fictitious magnetic flux of ±$\pm\pi/2$. Eq. (1) is invariant under T, which is an anti-unitary symmetry for spin-1/2. The eigenvalues of $\hat{W}_{p}$ change sign under T, thus forming Kramer’s pairs and rendering the overall spectrum two-fold degenerate. We henceforth denote the uniform $-\pi/2$ ($+\pi/2$) sector as the vortex-free (-filled) sector.
@@
$\hat{H}=\mathrm{i}\sum_{\langle j,k\rangle}J_{\alpha_{jk}}\hat{u}_{jk}\hat{c}_{j}\hat{c}_{k},$ (3)
where we introduced the Majorana bilinears $\hat{u}_{jk}=\mathrm{i}\hat{b}_{j}^{\alpha}\hat{b}_{k}^{\alpha}$. These link operators are anti-symmetric under link inversion $\hat{u}_{jk}=-\hat{u}_{kj}$ and under T, $\hat{u}_{jk}\overset{\mathrm{T}}{\longrightarrow}\hat{u}_{kj}$. The Majorana transformation effectively doubles the Hilbert space. Defining $\hat{D}_{i}=\mathrm{i}\hat{b}_{i}^{x}\hat{b}_{i}^{y}\hat{b}_{i}^{z}\hat{c}_{i}$ and $\hat{D}_{i}\ket{\psi}=+1\ket{\psi}\;\forall i$, the states {|$\{\ket{\psi}\}$ form the physical Majorana subspace.
Since each link couples only a one flavor of Majoranas, the operators $\hat{u}_{jk}$ commute with each other and Eq. (3). Therefore, we can replace them by the corresponding eigenvalues $u_{jk}=\pm 1$ relating to the vortex-free (filled) plaquette interpreted as a $\mathbb{Z}_{2}$ gauge field. Once a configuration of these eigenvalues is specified, the Hamiltonian (3) takes a non-interacting single-particle form. In order to express the gauge invariant eigenvalues $W_{p}$ of the Wilson loop operators by $u_{jk}$, we fix a direction for all links according to Fig. 1(b). Then, the Majorana transformation implies
$W_{p}=(-\mathrm{i})^{9}\prod_{\langle jk\rangle_{p}}u_{jk},$ (4)
where the product runs in a counterclockwise cycle over all links of a given plaquette.
@@
Figure 2: Energy cost of creating a vortex pair versus their distance in the hyperbolic plane on a droplet with 11878 sites. The energy cost increases with distance, i.e., indicating an attractive inter-vortex potential.
The energy of an eigenstate of (1) is given by the sum over the absolute value of all eigenvalues of (3) for a fixed plaquette flux configuration. For bipartite (and hence even-loop) $\{p,q\}$ lattices one can apply Lieb’s theorem [32] stating that the ground state lies in the vortex-free (vortex-filled) sector for $p\equiv 2\mod 4$ ($p\equiv 0\mod 4$). For the $\{9,3\}$ lattice, our numerical results indicate that the ground state(s) locate in the sectors with uniform vortex decoration. For the gapless domain parameter regime where $J_{\sigma(1)}<J_{\sigma(2)}+J_{\sigma(3)}\>\forall\sigma\{x,y,z\}$ with $\sigma$ denoting a 3-tuple permutation, we perform the following numerical experiments: (i) From the vortex-free sector, we insert a cluster of fluxes in a region of the hyperbolic lattice, and calculate the energy cost per vortex as well as the total energy difference to the vortex-free ground state (see Supplemental Material [33], Sec. A). The energy cost per new vortex is always positive, and the total energy cost increases with cluster size. (ii) We create a vortex pair and separate them from each other (Fig. 2), where we find an attractive inter-vortex potential. Based on these results, we conjecture that the vortex-free and vortex-filled sector ground state, which form a Kramers doublet, constitute the two-dimensional ground state manifold of (1).
Refer to caption Refer to caption (a)(b)
[…]
We calculate the bulk density of states of the Majorana Hamiltonian using the continued fraction method [34] (see [33], Sec. B). Representative points of parameter space are depicted in Fig. 3(a). Similar to the original Euclidean honeycomb lattice realization, we find two distinct regimes: a gapped phase $A$ and a gapless domain $B$ as depicted in Fig. 3(a). Furthermore, we find the phase boundary along the lines where one coupling constant is equal to the sum of the other two. Note that this symmetry of the phase diagram is linked to the canonical colorability of the $\{9,3\}$ lattice.
@@
The three regions $A_{x}$, $A_{y}$ and $A_{z}$, i.e., the dimerized phase regimes dominated by either strong Ising coupling, are equivalent modulo Ising axis rotation. Although the continued fraction method only converges slowly approaching the phase boundary to the gapless B domain, the band gap can be clearly identified. In order to further analyze the $A$ phases we consider the dimerized limit of strong coupling anisotropy $(J_{x},J_{y},J_{z})=(0,0,J)$ where the system reduces to decoupled Ising dimers. We contract the Ising dimers to a point, and regard them as effective spin-1/2 degrees of freedom on the $\{6,4\}$ hyperbolic lattice. There are fermionic and $\mathbb{Z}_{2}$ vortex excitations: A fermion is created when we break the alignment of two spins on a strong bond, which costs energy $\sim\!2J_{z}$. The low-energy physics of the $\mathbb{Z}_{2}$ vortices is governed by the effective Hamiltonian
$\hat{H}_{\text{eff}}^{(12)}=\text{const}$ $-\frac{143J_{x}^{4}J_{y}^{4}(J_{x}^{2}+J_{y}^{2})}{2^{17}J_{z}^{9}}\,\hat{W}_{p}\hat{W}_{p^{\prime}}$ (5)
$-\frac{91J_{x}^{6}J_{y}^{6}}{2^{20}J_{z}^{11}}\hat{W}_{p}\hat{W}_{p^{\prime\prime}},$
where $\hat{W}_{p}$ and $\hat{W}_{p^{\prime}}$ ($\hat{W}_{p^{\prime\prime}}$) are Wilson loop operators on two neighboring nonagons, sharing a weak (strong) bond (see [33], Sec. C). Flipping a bond $u_{jk}\to-u_{kj}$ creates a pair of vortex in the two adjacent nonagons. The resulting model is a hyperbolic version of the Kitaev Toric Code [35]. We do not further elaborate on the region $A$, as it largely corresponds to its Euclidean analogue. A subtle issue arises for topologically ordered ground states on hyperbolic lattices regarding the topological ground state degeneracy. The compactified hyperbolic space suggests a higher Riemann manifold genus which grows with system size. This implies a challenge in the reconciliation with the assumed relation between topological ground state degeneracy and genus [36], a principal issue which will be addressed elsewhere [37].
@@
The top left panels of Fig. 3(b) show the DOS in the gapless regime for representative points in parameter space. In contrast to the gapless regime for the Euclidean Kitaev model, our bulk spectra analysis indicates a finite, non-vanishing DOS at zero energy. This suggests a connection to the generically deviating character of hyperbolic semimetals [17], where the non-commutative geometry implies an effective Bloch band theory with higher-dimensional irreducible representations of translation. This deserves further investigation for the $\{9,3\}$ lattice in general. Moreover, the gapless phase we find here is intrinsically chiral since, due to the odd loops, it spontaneously breaks T symmetry in the thermodynamic limit. For small flakes up to $10^{5}$ sites, the single-particle Majorana representation of (1) can be diagonalized numerically [38]. In order to support our conjecture about the chiral character of phase $B$ we calculate the local Chern marker (LCM) proposed by Kitaev [7],

### v4

@@ Abstract
We analyze the Kitaev model on the $\{9,3\}$ hyperbolic lattice. The $\{9,3\}$ is formed by a regular tricoordinated tiling of nonagons, where the three-color coding of bonds according to the inequivalent Kitaev Ising spin couplings yields the natural generalization of the original Kitaev model for Euclidean regular honeycomb tiling. Upon investigation of the bulk spectrum for large finite size droplets, we identify a gapless chiral $\mathbb{Z}_{2}$ spin liquid state featuring spontaneous time-reversal symmetry breaking. Because of to its noncommutative translation group structure, such type of hyperbolic spin liquid is conjectured to feature chiral quasiparticles with a potentially non-Abelian Bloch profile.
@@
Figure 1: (a) Droplet of the $\{9,3\}$ hyperbolic lattice with Kitaev-colored bonds for each coupling $J_{x},J_{y},J_{z}$. The inset shows schematically the four Majorana fermions $\hat{b}_{j}^{\alpha},\hat{c}_{j}$ introduced at site $j$. (b) Direction convention for each link $\hat{u}_{jk}$. If all eigenvalues $u_{jk}$ are ±$\pm 1$, the ±$\pm\pi/2$ flux sector is realized. (c) Schematic phase diagram in the $J_{x}+J_{y}+J_{z}=1$ plane.
@@ Introduction.
Quantum spin liquids have been in the focus of contemporary research on strongly correlated electron systems for five decades [1]. In experiment, the unambiguous identification of such an intricate quantum fluid state of matter is notoriously difficult, and rarely, if ever, possible beyond reasonable doubt [2]. In theory, our understanding has progressed at various frontiers, yet at highly different rates. Due to the seminal work by Wen [3], an assumingly exhaustive classification of quantum spin liquids is given from their associated emergent gauge structure at low energies. There usually is, however, still a long way to go from quantum order classifications, and their related parton mean fields, to a refined microscopic realization and stabilization of such states. This is why, in particular for quantum spin liquids and quantum many-body phases in general, it is highly desirable to accomplish parent Hamiltonians whose ground state unambiguously matches such a state. Typically, a combination of optimization ansatz and educated guess is employed to scan a multidimensional Hamiltonian parameter space [4, 5, 6].
@@
The Kitaev paradigm of spin liquid parent Hamiltonians has proven particularly successful [7]. Conceived as a bond-dependent anisotropic Ising Hamiltonian between neighboring sites on the honeycomb lattice, it has brought about various exactly solvable instances of quantum spin liquids with an emergent $\mathbb{Z}_{2}$ gauge structure. In terms of spin liquids featuring topological order, it includes the gapped Abelian toric code liquid and, upon consideration of an external magnetic field [7] or through spontaneous time-reversal (T) symmetry breaking via lattice graphs with odd loops [8, 9], a gapped non-Abelian chiral Majorana liquid with spinon quasiparticles which resemble Pfaffian liquid quasiparticles [10, 11]. The Kitaev model further harbors the gapless $\mathbb{Z}_{2}$ spin liquid centered around isotropic Ising couplings. Further progress has been reached by Hermanns and Trebst [12] who first employed that if a unique bond attribution to three subsets can be made for a semiregular tricoordinated tiling, this still yields exact solvability of the Kitaev model. Upon transcending from two to three spatial dimensions, this enabled them to define Kitaev spin liquid states with a (potentially unstable [13]) spinon Fermi surface as well as chiral spin liquid states with crystalline gauge ordering of $\mathbb{Z}_{2}$ fluxes [14].
@@
Since the nature of the lattice strongly affects the magnetic frustration profile, any parent Hamiltonian search for quantum spin liquids seeks to consider all variations of potentially interesting lattices. Conventionally, this implies to be constrained to tessellations of Euclidean flat space. Recent progress on hyperbolic matter, however, has broadened the view to also consider Hamiltonians operating on negatively curved spaces and lattice tessellations thereof [15]. While the field initially started from unconventional noninteracting states and their description in terms of symmetry protected topological phases [16, 17, 18] and hyperbolic band theory [19, 20, 21, 22], it is now starting to reach the realms of quantum many-body physics through metal-insulator transitions, charge density wave states, and magnetism [23], fractional Chern insulators [24], and lattice-regularized formulations of AdS-CFT correspondence [25, 26]. An immediate platform for tabletop experimental realizations of hyperbolic space presents itself to be quantum and classical circuit networks [27, 28, 29].
@@
In this Letter, we introduce the notion of two-dimensional (2D) hyperbolic spin liquids in the framework of a generalized Kitaev model. Applying the same logical thread as in Euclidean space, the focus of consideration lies on tricoordinated lattice tessellations in the hope of reaching an exactly solvable Hamiltonian. While in principle any of the vast hyperbolic space tessellations that fulfill this tricoordination deserves further investigation, we will particularize to a microscopic model for an exotic spin liquid state that features gapless excitations, and yet spontaneously breaks time reversal symmetry. Notwithstanding previous conjectures about a possible chiral gapless spin liquid on the kagome lattice [30, 31], the microscopic setting and the exact ground state we find has no immediate analogue in Euclidean space, and hence expands the hitherto known landscape of spin liquids to uncharted territory.
@@
Realizing gapless excitations for a Kitaev Hamiltonian is intimately tied to regular tiling, i.e., tessellations with only one type of regular polygon, while spontaneous T symmetry breaking requires odd-sided loops. Such a tessellation of 2D Euclidean space by regular odd-sided tilings does not exist. Instead, one would have to resort to semi-regular tilings such as the Fisher star lattice [8, 32, 33] as well as non-Archimedean [9, 32] and amorphous [34] realizations, all of which feature gapped chiral Kitaev spin liquids. This motivates us to transcend to 2D hyperbolic space to realize tricoordinated, regular odd-loop tilings. Among the vast tessellations that fulfill the constraint, the $\{9,3\}$ lattice Fig. 1(a) stands out, as it features the smallest loop size of hyperbolic tiling whose bond coloring appears highly canonical, and even suggests itself to be the most natural generalization of the Euclidean honeycomb Kitaev model. Our analysis of the $\{9,3\}$ hyperbolic Kitaev model finds a gapless state which, due to the presence of odd loops, features spontaneous time-reversal symmetry breaking. We thus show the Kitaev model to harbor a gapless, chiral $\mathbb{Z}_{2}$ spin liquid state whose nature of excitations, however, is conjectured to be intimately dependent on the non-commutative translation group inherent to hyperbolic space.
Refer to caption Refer to caption Refer to caption (a)(b)(c)
[…]
@@
$\hat{H}=-J_{x}\!\sum_{\langle j,k\rangle_{x}}\hat{\sigma}_{j}^{x}\hat{\sigma}_{k}^{x}-J_{y}\!\sum_{\langle j,k\rangle_{y}}\hat{\sigma}_{j}^{y}\hat{\sigma}_{k}^{y}-J_{z}\!\sum_{\langle j,k\rangle_{z}}\hat{\sigma}_{j}^{z}\hat{\sigma}_{k}^{z}$ (1)
on a three-colored $\{9,3\}$ lattice. There, $\hat{\sigma}_{i}^{\alpha}$ denote Pauli matrices acting on site $i$, and $J_{\alpha}>0$ are the coupling constants along the respective link of type $\alpha=x,y,z$. The different link types are shown in Fig. 1(a) as red, green and blue lines, respectively. Although model (1) can be solved exactly for all tricoordinated three-colorable hyperbolic lattices, the $\{9,3\}$ appears to be the natural hyperbolic extension of the original Euclidean model due to the regular coloring scheme of repeated $x$, $y$, and $z$ links upon going counterclockwise around each plaquette (Fig. 1). Invoking $(p-2)(q-2)>4$ obeyed by all hyperbolic tilings, this canonical coloring scheme exists for all $\{p,3\}$ lattices with $p\equiv 0\mod 3,\;p>6$, so that the $\{9,3\}$ lattice is the one with smallest loop size.
The Hamiltonian (1) possesses a set of conserved quantities, one associated with each nonagon plaquette, given by the mutually commuting Wilson loop operators $\hat{W}_{p}=\hat{K}_{12}\hat{K}_{23}\hat{K}_{34}\hat{K}_{45}\hat{K}_{56}\hat{K}_{67}\hat{K}_{78}\hat{K}_{89}\hat{K}_{91}$, where $\hat{K}_{ij}=\hat{\sigma}_{i}^{\alpha}\hat{\sigma}_{j}^{\alpha}$, and the sites $i$ and $j$ are labeled by indices $i,j=1,\cdots,9$ counterclockwise around a plaquette and connected by a bond of type $\alpha$. Since all plaquette operators commute, the Hamiltonian attains a block diagonal form according to the $\hat{W}_{p}$ eigenvalues. In contrast to Kitaev models with even-loop plaquettes, these are given by ±$\pm\mathrm{i}$ rather than ±$\pm 1$, implying fictitious magnetic flux of ±$\pm\pi/2$. Equation (1) is invariant under $\mathrm{T}$, which is an antiunitary symmetry for spin $1/2$. The eigenvalues of $\hat{W}_{p}$ change sign under $\mathrm{T}$, thus forming Kramers pairs and rendering the overall spectrum twofold degenerate.
@@
$\hat{H}=\frac{\mathrm{i}}{2}\sum_{j,k}J_{\alpha_{jk}}\hat{u}_{jk}\hat{c}_{j}\hat{c}_{k},$ (2)
where we introduced the Majorana bilinears $\hat{u}_{jk}=\mathrm{i}\hat{b}_{j}^{\alpha}\hat{b}_{k}^{\alpha}$. These link operators are antisymmetric under link inversion $\hat{u}_{jk}=-\hat{u}_{kj}$ and under $\mathrm{T}$, $\hat{u}_{jk}\overset{\mathrm{T}}{\longrightarrow}\hat{u}_{kj}$. The Majorana transformation effectively doubles the Hilbert space. Defining $\hat{D}_{i}=\mathrm{i}\hat{b}_{i}^{x}\hat{b}_{i}^{y}\hat{b}_{i}^{z}\hat{c}_{i}$ and $\hat{D}_{i}\ket{\psi}=+1\ket{\psi}\;\forall i$, the states {|$\{\ket{\psi}\}$ form the physical Majorana subspace.
Since each link couples only one flavor of Majoranas, the operators $\hat{u}_{jk}$ commute with each other and Eq. (2). Therefore, we can replace them by the corresponding eigenvalues $u_{jk}=\pm 1$ relating to the ±$\pm\pi/2$ sectors interpreted as a $\mathbb{Z}_{2}$ gauge field. Once a configuration of these eigenvalues is specified, the Hamiltonian (2) takes a noninteracting single-particle form. In order to express the gauge invariant eigenvalues $W_{p}$ of the Wilson loop operators by $u_{jk}$, we fix a direction for all links according to Fig. 1(b). Then, the Majorana transformation implies $W_{p}=(-\mathrm{i})^{9}\prod_{\langle jk\rangle_{p}}u_{jk}$, where the product runs in a counterclockwise cycle over all links of a given plaquette $p$.
@@
Given a flux configuration, the energy of an eigenstate of (1) is computed by the sum over the absolute value of all eigenvalues of (2). For bipartite (and hence even-loop) $\{p,q\}$ lattices Lieb’s theorem [35] is applicable, stating that the ground state lies in the uniform flux-00 or flux-$\pi$ sector. For the $\{9,3\}$ lattice, a recent conjecture [34] as well as our numerical results (see Supplemental Material [36], Sec. A) indicate that the ground state(s) are characterized by uniform vortex decorations as well, such that the two-dimensional ground state manifold is constituted by the uniform ±$\pm\pi/2$ flux sectors, which form a Kramers doublet.
We calculate the bulk density of states (DOS) of the Majorana Hamiltonian using the continued fraction (CF) method [37] (see [36], Sec. B). Representative points of parameter space are depicted in Fig. 2(a). Although the CF expansion is not yet fully converged for the available system sizes, we identify two distinct regimes by the number of attractors of the CF coefficients: a gapped phase $A$ and a gapless domain $B$ (Fig. 2(b)). Furthermore, we find the phase boundary approximately along the lines where one coupling constant is equal to the sum of the other two. Note that the symmetry of the phase diagram is linked to the canonical colorability of the $\{9,3\}$ lattice.
@@
The three regions $A_{x}$, $A_{y}$ and $A_{z}$, i.e., the dimerized phase regimes dominated by either strong Ising coupling, are equivalent modulo Ising axis rotation. Although the CF method only converges slowly approaching the phase boundary to the gapless $B$ domain, the band gap can be clearly identified. In order to further analyze the $A$ phases we consider the dimerized limit of strong coupling anisotropy $(J_{x},J_{y},J_{z})=(0,0,J)$ where the system reduces to decoupled Ising dimers. We contract the Ising dimers to a point, and regard them as effective spin-1/2 degrees of freedom on the $\{6,4\}$ hyperbolic lattice. There are fermionic and $\mathbb{Z}_{2}$ vortex excitations: A fermion is created when we break the alignment of two spins on a strong bond, which costs energy $\sim\!2J_{z}$. The low-energy physics of the $\mathbb{Z}_{2}$ vortices is governed by the effective two-plaquette [38, 39] Hamiltonian
$\hat{H}_{\text{eff}}^{(10)}=$ $\,\text{const}+\frac{143\,J_{x}^{4}\,J_{y}^{4}}{2^{16}J_{z}^{9}}$
×$\times\bigg{(}J_{y}^{2}\sum_{\langle p,p^{\prime}\rangle_{x}}\hat{W}_{p}\hat{W}_{p^{\prime}}+J_{x}^{2}\sum_{\langle p,p^{\prime}\rangle_{y}}\hat{W}_{p}\hat{W}_{p^{\prime}}\bigg{)},$ (3)
where $\hat{W}_{p}$ and $\hat{W}_{p^{\prime}}$ are Wilson loop operators on two neighboring nonagons, sharing a weak $x$ or $y$ bond (see [36], Sec. C). Flipping a bond $u_{jk}\to-u_{jk}$ creates a pair of visons in the two adjacent nonagons, which increases the energy of the state by $\Delta E$ as the new $-\frac{\pi}{2}$ plaquettes are surrounded by $+\frac{\pi}{2}$ fluxes. Figure 3 compares $\Delta E$ observed by exact diagonalization of an OBC cluster with the result of Eq. (Dimerized phase A.). This effective model is a hyperbolic version of the Kitaev Toric Code [40]. We do not further elaborate on the region $A$, as it largely corresponds to its Euclidean analog. A subtle issue arises for topologically ordered ground states on hyperbolic lattices regarding the topological ground state degeneracy. The compactified hyperbolic space suggests a higher Riemann manifold genus which grows with system size. This implies a challenge in the reconciliation with the assumed relation between topological ground state degeneracy and genus [41], a principal issue which will be addressed elsewhere [42].
Figure 3: Energy increase $\Delta E$ due to vison pair creation separated by a $z$-type bond for $J_{x}=J_{y}=j\ll J_{z}$ in a 10 0241002410\,02410 024 site OBC cluster. The analytical value $\Delta E_{\textrm{lim}}$ predicted by Eq. (Dimerized phase A.) is recovered for $j\rightarrow 0$.
Refer to caption Refer to caption $t=0$Refer to caption $t=800$Refer to caption $t=1600$Refer to caption $t=2400$(a)(b)System size:10 0241002410\,02410 024 sites
Figure 4: (a) Phase diagram of the local Chern marker (LCM, cf. Eq. (4)) in the plane $J_{x}+J_{y}+J_{z}=1$ for a droplet with 1575 sites. The LCM is calculated from (2) including all states with negative energy. It is finite but not quantized in the chiral phase $B$, and vanishes in the $A$ phases. (b) Time evolution under (2) of a wave packet initially localized at the edge in the uniform $\pi/2$ flux sector for $J_{x}=J_{y}=J_{z}=\frac{1}{3}$. The center of the packet propagates unidirectionally along the boundary. The area of each circle is proportional to the probability to measure the excitation at the corresponding site.
@@
Panels (i), (ii) of Fig. 2(a) show the DOS in the gapless regime for representative points in parameter space. In contrast to the gapless regime for the Euclidean Kitaev model, our bulk spectra analysis indicates a finite, nonvanishing DOS at zero energy. This suggests a connection to the generically deviating character of hyperbolic semimetals [17], where the noncommutative geometry implies an effective Bloch band theory with higher-dimensional irreducible representations of translation. This deserves further investigation for the $\{9,3\}$ lattice in general. Moreover, the gapless phase we find here is intrinsically chiral since, due to the odd loops, it spontaneously breaks T symmetry in the thermodynamic limit. For flakes up to $10^{5}$ sites, the Majorana representation (2) can be diagonalized numerically [43]. To investigate the chiral character of phase $B$, we calculate the local Chern marker (LCM) proposed by Kitaev [7, 44],